Search
 
 

Display results as :
 


Rechercher Advanced Search

Latest topics
» Optimized heating profile for the CP, when CP and TM are 20cm distant, RH also around CP
Mon Mar 29, 2010 11:15 am by Alessio

» The SHACK-HARTMAN SENSOR as wavefront sensor
Thu Mar 18, 2010 12:41 pm by laura

» First test on the new CO2 laser
Tue Mar 16, 2010 6:35 pm by Alessio

» WI TCS beam alignment on ITM
Tue Mar 16, 2010 6:11 pm by Alessio

» Spot size change on NI photodiodes
Thu Mar 11, 2010 5:27 pm by laura

» TCS benches calibration
Thu Mar 11, 2010 5:24 pm by laura

» WI mirror B cabling
Thu Mar 11, 2010 5:20 pm by laura

» NI bench: power on the power meter was too low
Fri Feb 19, 2010 4:24 pm by Alessio

» NI TCS last activity
Fri Feb 19, 2010 4:23 pm by Alessio

Shopmotion


Affiliates
free forum
 


WI TCS beam alignment on ITM

Thu Mar 11, 2010 5:19 pm by laura

(Viviana, Alessio, Laura, F.Frasconi)

On tuesday afternoon we performed some operations required to allow closing of the WI tower.
1) placed a lens in front of the cross-hair to increase intensity of the arms of the cross.
2) aligned CO2 beam onto the cross-hair.
3) replaced hot mirror with 95%R beam splitter (to allow positioning of the IR camera, not yet done).
On wednesday morning, with …

[ Full reading ]

Comments: 1

Spot size change on NI photodiodes

Thu Mar 11, 2010 5:27 pm by laura

(Alessio, Viviana, Laura)

We notes that the size of the beam on the NI PDs is much smaller than that of the WI. So we moved the focus lens in front of the PDs in order to make the spot size approximately equal to that of the WI.

Here the distances between some optical components on the WI and NI respectively:

WI --> L-BS = 5cm ; BS-FOL = 12cm ; BS-FL = 7cm
NI --> L-BS = 11cm ; BS-FOL …

[ Full reading ]

Comments: 0

TCS benches calibration

Thu Mar 11, 2010 5:24 pm by laura

(Viviana, Alessio, Laura)

Last week, we also preformed the calibration of the TCS benches.

Figure 1 shows the NI TCS power versus the wave plate angle. Second plot shows the CO2 power as a function of the ADV voltage.
Minimum power is about 50mW, maxumim is 9.2W.

Calibration factor is:
PNI(W)=4.814*TCS_NI_Power_DC - 0.117

Figures 3 and 4 represent the same quantities for the WI. Minimum …

[ Full reading ]

Comments: 0

WI mirror B cabling

Thu Mar 11, 2010 5:20 pm by laura

(Viviana, Alessio, Laura, P.Ruggi, F. Gherardini)

On thursday morning, we re-connected the cables of the stepping motors to mirror B and checked that all the motors are working and the mirror is moving (1000 steps equals approximately 3cm on TM). We checked the end-stops of the motors.


[ Full reading ]

Comments: 0

NI bench: power on the power meter was too low

Fri Feb 19, 2010 4:24 pm by Alessio

This afternoon the opertator (see Operation report - Afternoon shift) reported a warning from ALP: the NI laser possibly being off due to low signal on TCS_power_DC.

After a quick investigation, I saw that the power on the power meter
was very low because of the setting of the wave-plate angle. So I
simply rotated the plate to shine more power on the power meter: the
flag on the DMS is green …

[ Full reading ]

Comments: 0

NI TCS last activity

Fri Feb 19, 2010 4:23 pm by Alessio

(Alessio, Viviana, F. Frasconi, F. Gherardini, P. Ruggi)

Today, we finished to fix the mirror B holder.
Then we re-connected the cables of the stepping motors to mirror
B and checked that all the motors are working and the mirror is moving.
Finally we also checked the end-stops of the motors. The screen used
to position the cross-hair on the ITM is still in place.
The re-installation of the …

[ Full reading ]

Comments: 0

NI TCS beam alignment on ITM

Fri Feb 19, 2010 4:22 pm by Alessio

(Alessio, Viviana, P. Puppo. F. Frasconi)
Yesterday we performed the alignment of the TCS beam on the optics inside the NI tower.
With Franco, we re-installed mirror B and removed the protection from mirror A.
Then, with Franco and Paola, we looked for any diffusion of the
cross-hair on the first contact film of the ITM and nothing was
absolutely visible. We tried to put a diffusive screen …

[ Full reading ]

Comments: 0

IR camera results investigation

Fri Feb 12, 2010 5:06 pm by Alessio

(Alessio, J. Marque, Viviana)

As reported in Last NI bench adjustment,
after the installation of the IR camera in its design position, we
noticed a discrepancy between the IR camera image and the usual thermal
image.
Figure 1 shows the thermal camera image and figure 2 shows the
corresponding IR camera image. It is possible to see that the ring
looks fine on the thermal camera and rather …

[ Full reading ]

Comments: 0

Last NI bench adjustment

Fri Feb 12, 2010 5:03 pm by Alessio

(Alessio, Laura, Viviana)

Yesterday we performed some operations required to allow closing of the NI tower.
1) placed a lens in front of the cross-hair to reduce size and to increase intensity of the arms of the cross.
2) carefully aligned CO2 beam onto the cross-hair.
3) replaced hot mirror with 95%R beam splitter and positioned the IR camera in their design position.
During last operation we …

[ Full reading ]

Comments: 0

Statistics
We have 8 registered users
The newest registered user is laura

Our users have posted a total of 287 messages in 160 subjects
Who is online?
In total there is 1 user online :: 0 Registered, 0 Hidden and 1 Guest

None

[ View the whole list ]


Most users ever online was 17 on Wed Nov 21, 2012 7:34 am